• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 18:42:46 on March 9,2002

    Entry number 1280

    keyword=Shift Summary

    16:00 started BCM calibration; first did inject then hall, inject then hall to get the hall bcm linearity. Followed by injector (all currents) and hall (all currents) to calibrate the bcms.

    18:00 BCM finished. Inserted BeO to check spot location. Target camera does not show up. Moving to 15 cm target

    Set raster to 2x2 mm iteratively (it seems x and y maybe swapped between our settings and MCC settings). MCC has it set to 1.5 x 2 mm.

    18:45 Analyzed run 4051 (40 uA) to be sure detectors work and target is OK. Everything looks good so we are going to 70 uA, then 100 uA.